OpenStudy (calculusxy):

If you have six people in a room, and each person shakes hands with every other person exactly once, how many total handshakes happen? @phi @amistre64 @AkashdeepDeb @hartnn @ranga

OpenStudy (anonymous):

36

OpenStudy (calculusxy):

There are no choices of 36.

OpenStudy (anonymous):

what are the choices?

OpenStudy (akashdeepdeb):

From 6 people you have TO SELECT 2 poeple for a hand shake. So the solution would be 6C2 = 6 * 5 / 2 * 1 = 15 ways! :D

OpenStudy (anonymous):

you wouldn't know because they shake hands with every other person?

OpenStudy (calculusxy):

@AkashdeepDeb Can you give me a more easier way to solve this problem?

OpenStudy (akashdeepdeb):

Do you know combinations? :)

OpenStudy (calculusxy):

No

OpenStudy (calculusxy):

Can u explain it to me :)

OpenStudy (akashdeepdeb):

Sure, but first, Do you know factorials?

OpenStudy (calculusxy):

No

OpenStudy (akashdeepdeb):

I think then it'll take you some time and wouldn't be possible to teach on OS.

OpenStudy (phi):

one way is say: one person will shake hands with 5 other people there are 6 people so 5*6 = 30 but, we are double counting (A with B is the same as B with A) so 30/2 = 15 handshakes.

OpenStudy (phi):

but 6 choose 2 is a nicer way to solve this problem

OpenStudy (akashdeepdeb):

@phi 's explanation Is TOTALLY CORRECT. That IS in fact, in short, the function of combinations in probability and in the fundamental principal of counting.

OpenStudy (calculusxy):

Thanks @AkashdeepDeb and @phi. I learned something new today :)

OpenStudy (akashdeepdeb):

:)

hartnn (hartnn):

one more way to look at it is, the first person will shake hands with 5 other person, so \(\huge5\) the 2nd person already shaked his hand with the 1st person, remaining shakes = \(\huge4\) continuing this way, we get 5+4+3+2+1 = 15 ways infact there a formula dedicated to this handshake problem, which directly comes from sum of 'n' natural number formula, for n ppl, total handshakes will be \(\large \dfrac{n(n-1)}{2}\)